The profit on a watch is given by P = x^2 – 5x – 9 and where x is the number of watches sold per day. How many watches were sold on a day when there was a $15 loss?
Answer by nerdybill(5407)

The profit on a watch is given by P = x^2 – 5x – 9 and where x is the number of watches sold per day. How many watches were sold on a day when there was a $15 loss?
.
Simply set P to -15 and solve for x:
P = x^2 – 5x – 9
15 = x^2 – 5x – 9
0 = x^2 – 5x – 24
0 = (x-8)(x+3)
x = {-3, 8}
.
Toss out the negative solution leaving:
x = 8